OpenStudy (anonymous):

What is the distance, to the nearest hundredth, from (6, 5) to (-3, -8)? Select one: a. 4.24 b. 16.12 c. 15.81 d. 250

OpenStudy (anonymous):

You will need to use the distance formula \[\sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\] So you will have \[\sqrt{(-3-6)^{2}+(-8-5)^{2}}\] \[\sqrt{81+169}\]

OpenStudy (anonymous):

so its 15.81

OpenStudy (anonymous):

Should be